[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Oct 21, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Fri Oct 21 13:06:59 UTC 2016


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
no new problems
HSU:
no new problems
Other items:
Forecasters reported a windows problem going to certain webpages such as  
twitter. OPC also reported a problem. NCO is working on a fix.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161021/97f0f653/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list